# Symmetry Breaking in Chemical Systems: Engineering Complexity through Self-Organization and Marangoni Flows

### v3

@@ 4 Method Section
The following chemicals and materials were used in this study. Sodium Bromate (NaBrO3), Sodium Bromide (NaBr), Conc. Sulfuric acid (Conc. H2SO4), 1,4-Cyclohexanedione, Ferrous sulfate, 1,10-phenanthroline was provided by sigma and used without any further purification. Ferroin was synthesized by mixing Ferrous sulfate hexahydrate and 1,10-phenanthroline in 1:3 molar ratio. For CHD-BZ experiments different concentration of the stock solutions were prepared in dH2O. Solution A: (494.58 mM solution of sodium bromate) Solution B: (891.9 mM solution of 1,4-Cyclohexanedione) Solution C: (971.9 mM solution of NaBr). A mixture was prepared in a beaker by combining 9 mL of solution A with 1.5 mL of solution B, which was then homogenized using a magnetic stirrer. Following this, 1.5 mL of the Ferroin indicator was added to the mix. Under a chemical fume hood, 0.8 mL of solution C was cautiously introduced to the reaction. The stirring continued for an estimated 1 hour and 45 minutes. Subsequent to this phase, an additional 1.5 mL of Ferroin was incorporated into the mixture. The stirring persisted at a controlled temperature of 23°C for a further 30 minutes, after which it was halted. Attention was then given to the mixture for any signs of color transformation, signaling the start of wave propagation, typically observed 3.5 to 4 hours post the initial mixing process. Following the onset of color change, the CHD-BZ reaction mixture was transferred to a plasma-treated PDMS mold featuring an array of macro-pillars, as specified in Figure 1A. These pillars, typically 1 mm in diameter and 3 mm high, are arrayed in a square grid with 5 mm spacing. In experiments with variable pillar dimensions and spacings, the volume of the CHD-BZ solution was adjusted to ensure a liquid height of about 1.5 mm, thus avoiding submergence of the pillars. Upon pouring, the solution was swirled gently until the mixture achieved a brown color. The Petri dish was then positioned undisturbed beneath a camera to record the wave propagation, under the illumination of white LED light. For each experiment, we record the waves for 5–10 minutes, observing until the waves from the Petri dish boundary interact with those centered around the obstacles. Afterward, we stop the recording, swirl the solution in the Petri dish to reset the system, and start recording again. Typically, the reagents are fully depleted after about 2 hours. In some experiments, a smaller, modified 50 mm Petri dish was used, and the volume of the CHD-BZ solution was correspondingly reduced to 1.5 mL to maintain the same liquid height. It’s important to note that in our standard setups, unless specified otherwise, the pillars do not submerge in the chemical solution. The hydrophilic nature of plasma-treated PDMS causes a thin layer of liquid to ascend the pillars (Figure 1B), altering the liquid surface near the obstacles as depicted in Figure S2C.
For some control experiments, the surface of the PDMS was coated with a 20 nm titanium layer and then with a pure gold layer. This was done using a Physical Vapor Deposition (PVD) technique with the Lesker Proline PVD 75 apparatus, rotating the substrate four times to achieve uniform coverage on the PDMS pillars’ surfaces.
@@
Image Processing
The wave period is roughly 45 sec, and we capture images of the chemical waves at a rate of 1 frame per second across multiple wave cycles. To minimize spatial noise, we average the captured images over several wave cycles, typically 6-8, and then subtract each individual image from the time-averaged image. After this, we apply a Gaussian blur filter and use a customized MATLAB code to perform the Hilbert transform, which helps us calculate the phase map. This code is provided in the Supplementary Information.
@@
Numerical simulations
We used the three-component Tyson-Fife model [80] to simulate our experimental system. The model focuses on the essential dynamics of the reaction by reducing the full chemical mechanism to a manageable number of reactions and variables. It typically includes the inhibitor $w$ which changes on the fastest time scale, activator $u$ which is the second fastest variable, and the slow varying inhibitor $v$. The variables $u$, $v$ and $w$ represent the concentrations of bromous acid (HBrO2), an oxidized form of the catalyst (such as ferroin Fe3+), and bromide ions, respectively. The set of equations for activator $u$ and two inhibitors $v$ and $w$ are:
$\frac{du}{dt}$ $=$ $D\nabla^{2}u+\frac{1}{\epsilon}(aw-uw+u-u^{2}),$ (1)
$\frac{dv}{dt}$ $=$ $u-v,$ (2)
$\frac{dw}{dt}$ $=$ $\frac{1}{\epsilon^{\prime}}(-aw-uw+bv),$ (3)
where $\epsilon^{\prime}\ll\epsilon\ll 1$, $a\ll 1$ and $b\simeq 1$. The parameters used in our simulations are based on those in Ref. [80]: $D=1$, ϵ=italic-$\epsilon=$0.05, ϵ=superscriptitalic-$\epsilon^{\prime}=$0.001, $b=$2.5, $a=$0.002 . We note that at $b=$2.5, the system is in the excitable state. The dimensions of the simulation box are chosen to be $N_{x}=N_{y}=1023$. The positions of the pillars on a square lattice are determined by $x$ and $y$ coordinates that meet the condition:$0.5[1+\cos(k(x-y)\cos(\pi/4))*\cos(k(x+y)\cos(\pi/4))]\leq 0.1$, where $k^{-1}$ representing the spacing between the pillars, is chosen to be $N_{x}/(6\pi)$. The concentration of inhibitor $w$ is set to zero inside the pillars.
In the simulations featuring an asymmetric distribution of $w$ around the pillars, we model the concentration of $w$ in a narrow ring around the pillars as: $w(x,y)=\alpha[1+\cos(k^{\prime}(x-y)\cos(\pi/4))*\cos(k^{\prime}(x+y)\cos(\pi/4))]$, where $k^{\prime}$ is arbitrarily set to $\pi/10$ and $\alpha=0.5$. After one cycle of circular waves, the wavefront breaks down and produces four petals, with the size of each petal varying over time, as depicted in Figure 12C.
In the tilted Petri dish experiment shown in Figure S5A-B, we adjusted the model to include a term proportional to the height $h(x)=x/N_{x}$ which varies linearly with $x$, under the assumption that the fluid height changes solely in the $x$-direction:
$\frac{du}{dt}$ $=$ $D\nabla^{2}u+\frac{1}{\epsilon}(aw-uw+u-u^{2}+p_{1}h(x)),$ (4)
$\frac{dv}{dt}$ $=$ $u-v+p_{2}h(x),$ (5)
$\frac{dw}{dt}$ $=$ $\frac{1}{\epsilon^{\prime}}(-aw-uw+bv+p_{3}h(x)),$ (6)
where $p_{1}=p_{2}=p_{3}=0.01$.
@@ Acknowledgements
We deeply appreciate the valuable insights and discussions provided by Professors G. Shubeita, A. Narayanan, A. Rebane, A. Pumir, O. Omelchenko, E. Frey, and D. Lohse. We also thank Dr. I. Gholami and M. Ibrahim for technical assistance and to the anonymous reviewers for their thoughtful comments and suggestions. We are especially grateful to the Core Technology Platform (CTP) at NYUAD for their exceptional support, with special thanks to Dr. R. Rezgui for his critical assistance with microscopy, Dr. Q. Zhang for his support in the clean room, and V. Dhanvi and J. Govindan for their outstanding contributions to fabricating the acrylic molds. Finally, we acknowledge the High Performance Computing (HPC) facility at NYUAD for providing essential resources for data storage and analysis, which were integral to this study.
Author contributions: S.G. and B.P. conducted the experiments, O.S. and A.G. designed and carried out the numerical simulations. A.G. conceived the idea, designed the experiments, analyzed the data, and wrote the first draft of the manuscript. All authors discussed the results and revised the manuscript.
Data and materials availability All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. Additional data related to this paper may be requested from the authors.
S5 Supplemental Figures
@@
Figure S5: Phase map of an experiment in a tilted Petri dish (A) In a Petri dish with a slight tilt, synchronization waves initiate in the thinner part of the fluid and propagate towards the area with greater fluid thickness (video 15). (B) Space-time plot of the phase map along the green dashed line indicated in panel (A) illustrates the propagation of phase waves. Notice that in this experiment, the Petri dish is not covered, and flower patterns emerge at $t\sim 80$ sec. (C) Numerical simulations emulating the experimental setup shown in panel (A).
Figure S6: Numerical simulations with asymmetric $w$ around the pillar. (A) The space-time plot along the line marked in panel (B) illustrates that the obstacles become wave centers. (B) shows an image of the circular waves emanating from the obstacles. (C-D) Over time, the wavefront breaks, leading to the emergence of flower-like wave patterns.
Figure S7: Marangoni-driven surface flows. (A-B) In this experiment, tracer particles (20 $\mu$m in diameter) are used to visualize surface flows driven by Marangoni flows. These flows arise from chemical concentration gradients and slight temperature differences caused by evaporative cooling, which generate surface tension gradients (refer to video 27).

### v4

@@ 4 Method Section
Chemical preparation The following chemicals and materials were used in this study. Sodium Bromate (NaBrO3), Sodium Bromide (NaBr), Conc. Sulfuric acid (Conc. H2SO4), 1,4-Cyclohexanedione, Ferrous sulfate, 1,10-phenanthroline was provided by sigma and used without any further purification. Ferroin was synthesized by mixing Ferrous sulfate hexahydrate and 1,10-phenanthroline in 1:3 molar ratio. For CHD-BZ experiments different concentration of the stock solutions were prepared in dH2O. Solution A: (494.58 mM solution of sodium bromate) Solution B: (891.9 mM solution of 1,4-Cyclohexanedione) Solution C: (971.9 mM solution of NaBr). A mixture was prepared in a beaker by combining 9 mL of solution A with 1.5 mL of solution B, which was then homogenized using a magnetic stirrer. Following this, 1.5 mL of the Ferroin indicator was added to the mix. Under a chemical fume hood, 0.8 mL of solution C was cautiously introduced to the reaction. The stirring continued for an estimated 1 hour and 45 minutes. Subsequent to this phase, an additional 1.5 mL of Ferroin was incorporated into the mixture. The stirring persisted at a controlled temperature of 23°C for a further 30 minutes, after which it was halted. Attention was then given to the mixture for any signs of color transformation, signaling the start of wave propagation, typically observed 3.5 to 4 hours post the initial mixing process. Following the onset of color change, the CHD-BZ reaction mixture was transferred to a plasma-treated PDMS mold featuring an array of macro-pillars, as specified in Figure 1A. These pillars, typically 1 mm in diameter and 3 mm high, are arrayed in a square grid with 5 mm spacing. In experiments with variable pillar dimensions and spacings, the volume of the CHD-BZ solution was adjusted to ensure a liquid height of about 1.5 mm, thus avoiding submergence of the pillars. Upon pouring, the solution was swirled gently until the mixture achieved a brown color. The Petri dish was then positioned undisturbed beneath a camera to record the wave propagation, under the illumination of white LED light. For each experiment, we record the waves for 5–10 minutes, observing until the waves from the Petri dish boundary interact with those centered around the obstacles. Afterward, we stop the recording, swirl the solution in the Petri dish to reset the system, and start recording again. Typically, the reagents are fully depleted after about 2 hours. In some experiments, a smaller, modified 50 mm Petri dish was used, and the volume of the CHD-BZ solution was correspondingly reduced to 1.5 mL to maintain the same liquid height. It’s important to note that in our standard setups, unless specified otherwise, the pillars do not submerge in the chemical solution. The hydrophilic nature of plasma-treated PDMS causes a thin layer of liquid to ascend the pillars (Figure 1B), altering the liquid surface near the obstacles as depicted in Figure S2C. For some control experiments, the surface of the PDMS was coated with a 20 nm titanium layer and then with a pure gold layer. This was done using a Physical Vapor Deposition (PVD) technique with the Lesker Proline PVD 75 apparatus, rotating the substrate four times to achieve uniform coverage on the PDMS pillars’ surfaces.
@@
Image Processing The wave period is roughly 45 sec, and we capture images of the chemical waves at a rate of 1 frame per second across multiple wave cycles. To minimize spatial noise, we average the captured images over several wave cycles, typically 6-8, and then subtract each individual image from the time-averaged image. After this, we apply a Gaussian blur filter and use a customized MATLAB code to perform the Hilbert transform, which helps us calculate the phase map. This code is provided in the Supplementary Information.
@@
Numerical Simulations Our numerical simulations qualitatively capture the hydrodynamic instabilities observed in CHD-BZ experiments. Our aim is to demonstrate that solutal Marangoni–driven instabilities, induced by evaporation, suffice to explain the symmetry-breaking patterns seen experimentally. We do not attempt a quantitative match to the data, as that would require a substantially more complex model.
We model the system using the experimental geometry: a binary mixture exhibiting Marangoni instabilities – one example being an ethanol–water system – is confined around a cylindrical pillar of radius $R=$1\text{\,}\mathrm{mm}$\$, height of $5\text{\,}\mathrm{mm}$, and center-to-center spacing of $7\text{\,}\mathrm{mm}$ (see Figure 12). At the evaluation time $t=599$ s—recorded from our experiments—the liquid layer meets the pillar walls at a contact angle $\theta=52^{\circ}$ and has a height of $2.5\text{\,}\mathrm{mm}$ there. The evaporation is assumed to be limited by vapor diffusion, a reasonable assumption for liquids sufficiently below the boiling point. The contact line remains pinned to the pillar walls, and a symmetry boundary is imposed at the midpoint between adjacent pillars. Evaporation is modeled using the quasi-stationary framework of Rocha et al. [83], suitably adapted to this geometry.
We track the mass fraction $y_{A}$ of the more volatile component $A$ (e.g., ethanol), with $y_{B}=1-y_{A}$ denoting the mass fraction of the less volatile component $B$ (e.g., water). For slowly evaporating liquids, compositional variations remain small, so we decompose
$y_{A}(x,t)=y_{A,0}(t_{0})+y(x,t)\,$, (1)
where $y_{A,0}(t_{0})$ is the spatially averaged mass fraction at the evaluation time $t_{0}$ and $y(x,t)$ captures local deviations.
Although $y_{A,0}(t)$ evolves over time, the flow can be treated as quasi‐steady at each drying stage, provided the evaporation rate remains low. All material properties are held constant except for surface tension, which is a function of the mass fraction variations. We assume a linear dependence
$\sigma(y_{A})=\sigma\bigl{(}y_{A,0}(t)\bigr{)}+y\,\frac{\partial\sigma}{\partial y_{A}}\,.$
The system is nondimensionalized as follows:
$x=R\,\tilde{x},\quad t=\frac{R^{2}}{D_{0}}\,\tilde{t},\quad\mathbf{u}=\frac{D_{0}}{R}\,\tilde{\mathbf{u}},\quad p=\frac{\mu_{0}\,D_{0}}{R^{2}}\,\tilde{p},\quad c_{\alpha}-c_{\alpha}^{\infty}=\bigl{(}c_{\alpha}^{\text{eq}}-c_{\alpha}^{\infty}\bigr{)}\,\tilde{c}_{\alpha},$
where $D_{0}$ and $\mu_{0}$ are the diffusion coefficient and viscosity at the mean composition,
$c_{\alpha}^{\text{eq}}=c_{\alpha}^{\text{pure}}\,\gamma_{\alpha}\,x_{\alpha}$
is the equilibrium vapor concentration (with $c_{\alpha}^{\text{pure}}$ the pure‐component saturation, $\gamma_{\alpha}$ the activity coefficient, and $x_{\alpha}$ the mole fraction), and $c_{\alpha}^{\infty}$ is the far‐field vapor concentration.
Assuming small local variations in $y$, we treat $c_{\alpha}^{\text{eq}}$ as constant [80]. Under this scaling, the dimensionless vapor field $\tilde{c}$ in the gas phase satisfies the Laplace equation
$\tilde{\nabla}^{2}\tilde{c}=0$ (2)
subject to the boundary conditions $\tilde{c}=0$ at the liquid–vapor interface and $\tilde{c}=1$ in the far field [89, 90, 80]. The solution $\tilde{c}$ then maps directly back to the physical vapor concentration $c_{\alpha}$. The dimensionless evaporation rate $\tilde{j}$ is given by the diffusive mass flux
$\tilde{j}=-\mathbf{n}\boldsymbol{\cdot}\tilde{\boldsymbol{\nabla}}\tilde{c}$, (3)
with $\mathbf{n}$ being the outward normal vector at the liquid-vapour interface. Evaporation drives variations in the liquid composition through the flux boundary condition:
$\tilde{\boldsymbol{\nabla}}y\boldsymbol{\cdot}\mathbf{n}=\mathrm{Ev}\tilde{j}$, (4)
where $\mathrm{Ev}$ is the evaporation number, which was originally defined by [80] as
$\mathrm{Ev}=\dfrac{(1-y_{A,0})D_{A}^{\mathrm{vap}}(c_{A,0}^{\mathrm{eq}}-c_{A}^{\infty})-y_{A,0}D_{B}^{\mathrm{vap}}(c_{B,0}^{\mathrm{eq}}-c_{B}^{\infty})}{\rho_{0}D_{0}},$ (5)
where $D_{A}^{\mathrm{vap}}$ and $D_{B}^{\mathrm{vap}}$ are the vapour diffusion coefficients of components $A$ and $B$, respectively, and $\rho_{0}$ is the liquid density evaluated at the spatially averaged composition.
In the liquid phase, we consider the incompressible Stokes equations (neglecting inertia and gravity):
$\tilde{\boldsymbol{\nabla}}\boldsymbol{\cdot}\tilde{\mathbf{u}}=0,$ (6)
$\tilde{\boldsymbol{\nabla}}p=\tilde{\boldsymbol{\nabla}}^{2}\tilde{\mathbf{u}}$. (7)
In order to reduce the number of parameters in the problem, we define $\xi=y/\mathrm{Ev}$, such that Eq. (4) becomes:
$\tilde{\boldsymbol{\nabla}}\xi\boldsymbol{\cdot}\mathbf{n}=\tilde{j}$. (8)
The evolution of $\xi$ in the liquid phase is then given by the advection-diffusion equation:
$\partial_{\tilde{t}}\xi+\tilde{\mathbf{u}}\boldsymbol{\cdot}\tilde{\boldsymbol{\nabla}}\xi=\tilde{\nabla}^{2}\xi.$ (9)
Marangoni stresses are considered at the liquid-vapour interface via the boundary condition:
$\mathbf{n}\boldsymbol{\cdot}(\tilde{\boldsymbol{\nabla}}\tilde{\mathbf{u}}+(\tilde{\boldsymbol{\nabla}}\tilde{\mathbf{u}})^{t})\boldsymbol{\cdot}\mathbf{t}=\mathrm{Ma}\tilde{\boldsymbol{\nabla}}_{t}\xi\boldsymbol{\cdot}\mathbf{t},$ (10)
with $\mathbf{t}$ being the tangential vector at the liquid-vapour interface and $\mathrm{Ma}$ being the solutal Marangoni number, defined as
$\mathrm{Ma}=\frac{R\partial_{y_{A}}\sigma}{\mu_{0}D_{0}}\mathrm{Ev}.$ (11)
We consider that the shape of the liquid layer does not change due to surface tension effects (i.e. capillary number $\mathrm{Ca}\rightarrow 0$). At the liquid-gas interface, we consider that evaporation is slow, i.e. $\mathbf{u}_{I}\ll\mathbf{u}$, where $\mathbf{u}_{I}$ is the velocity at the interface and $\mathbf{u}$ is the velocity in the bulk liquid. This allows us to assume that the interface is stationary, i.e. we impose the boundary condition $\tilde{\mathbf{u}}\boldsymbol{\cdot}\mathbf{n}=0$ at the liquid-vapour interface, which we enforce via a locally adjusted normal traction. The volume of liquid in the system is fixed to a value such that the contact angle is $\theta=52^{\circ}$ at the pillar walls and the liquid is pinned to the pillar walls. At the substrate, we impose a no-slip boundary condition $\tilde{\mathbf{u}}=0$ and a no-penetration boundary condition $\boldsymbol{\nabla}\xi\boldsymbol{\cdot}\mathbf{n}=0$. We constraint the spatial average of $\xi$ to zero, in accordance with the decomposition (1).
The system of equations is solved using the finite element method with the open-source library pyoomph [91], which is based on oomph-lib [92] and GiNaC [93]. All domains are discretised using triangular elements. The operators in the system of equations are given in a cylindrical coordinate system. Linear basis functions are used for $\xi$ and $\tilde{p}$, whilst quadratic basis functions are used for the velocity field $\tilde{\mathbf{u}}$. We follow the methodology outlined in [83] to compute the steady base state of the system and its stability against perturbations characterised by different azimuthal wavenumbers $m$. This allows us to assess whether the system’s symmetry is broken and forms patterns with a specific azimuthal wavenumber, like those observed in the experiments.
The stability of the simplified system depends on the Marangoni number $\mathrm{Ma}$, on the distance between pillars $D$, on the angle $\theta$ at the pillar walls, and on the liquid height at the pillar walls. In the following, for simplicity, we fix the liquid height at $2.5$ mm, $D=7$ mm, $\theta=52^{\circ}$, and vary $\mathrm{Ma}$.
@@ Acknowledgements
We deeply appreciate the valuable insights and discussions provided by Professors G. Shubeita, A. Narayanan, A. Rebane, A. Pumir, O. Omelchenko and E. Frey. We also thank Dr. I. Gholami and M. Ibrahim for technical assistance and to the anonymous reviewers for their thoughtful comments and suggestions. We are especially grateful to the Core Technology Platform (CTP) at NYUAD for their exceptional support, with special thanks to Dr. R. Rezgui for his critical assistance with microscopy, Dr. Q. Zhang for his support in the clean room, and V. Dhanvi and J. Govindan for their outstanding contributions to fabricating the acrylic molds. We also acknowledge the High Performance Computing (HPC) facility at NYUAD for providing essential resources for data storage and analysis, which were integral to this study. Finally, this work was also supported by an Industrial Partnership Programme, High Tech Systems and Materials (HTSM), of the Netherlands Organisation for Scientific Research (NWO); a funding for public-private partnerships (PPS) of the Netherlands Enterprise Agency (RVO) and the Ministry of Economic Affairs (EZ); Canon Production Printing Netherlands B.V.; University of Twente; and Eindhoven University of Technology (project TKI HTSM - CANON - P1 - PRINTHEAD & DROPLET FORMATION, grant no. PPS2107).
Author contributions
S.G. and B.P. conducted the experiments, D.R. , C.D. and D.L. designed and carried out the numerical simulations. A.G. conceived the idea, designed the experiments, analyzed the experimental data, and wrote the first draft of the manuscript. All authors discussed the results and revised the manuscript.
Data and materials availability
All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. Additional data related to this paper may be requested from the authors.
Supplemental Figures
@@
Figure S5: Phase map of an experiment in a tilted Petri dish (A) In a Petri dish with a slight tilt, synchronization waves initiate in the thinner part of the fluid and propagate towards the area with greater fluid thickness (video 15). (B) Space-time plot of the phase map along the green dashed line indicated in panel (A) illustrates the propagation of phase waves. Notice that in this experiment, the Petri dish is not covered, and flower patterns emerge at $t\sim 80$ sec.
[…]
@@
Meniscus analysis
At static equilibrium, the pressure jump across a liquid–air interface equals surface tension times mean curvature (Young–Laplace law):
$\gamma\,\kappa\;=\;\Delta P.$ (S.1)
Measuring height $H(r)$ upward (relative to a far-field reference) gives the hydrostatic pressure difference $\Delta p=\rho g\,H$, where $\rho$ is the liquid–air density difference and $g$ is gravity. For an axisymmetric meniscus $z=H(r)$ around a vertical cylinder, the exact mean curvature is the sum of principal curvatures:
$\kappa\;=\;\,\frac{H^{\prime\prime}}{\bigl{(}1+H^{\prime 2}\bigr{)}^{3/2}}\;+\;\frac{1}{r}\,\frac{H^{\prime}}{\sqrt{1+H^{\prime 2}}},$ (S.2)
where primes denote derivatives with respect to $r$.
Figure S9: Meniscus extraction, model fit, and pressure reconstruction. (A) Side-view of the rising film around a cylindrical obstacle; the interface (yellow) is obtained by edge tracing within a user-defined region of interest, and the dashed line marks the cylinder wall. (B) Dimensionless meniscus profile (circles) together with the best-fit far-field model that enforces the wall contact angle; the example fit returns lambda 1.7, offset -0.0485, and a fitted Bond number of 2.91. (C) Laplace pressure reconstructed along the interface, shown in both dimensionless and dimensional units; the solid red curve uses the exact curvature while the dashed black curve shows a linearized approximation, with the largest differences near the wall.
Let $R$ be the cylinder radius and define nondimensional variables $\tilde{r}=r/R$ and $\tilde{H}=H/R$. Combining (S.1)–(S.2) yields the exact (nonlinear) shape equation
$\frac{d}{d\tilde{r}}\left[\frac{\tilde{r}\,\tilde{H}^{\prime}}{\sqrt{1+(\tilde{H}^{\prime})^{2}}}\right]\;-\;\mathrm{Bo}\,\tilde{r}\,\tilde{H}\;=\;0,\qquad\mathrm{Bo}\equiv\frac{\rho gR^{2}}{\gamma},$ (S.3)
with boundary conditions
$\tilde{H}^{\prime}(1)\;=\;-\cot\theta,\qquad\tilde{H}(\tilde{r})\to 0\quad\text{as}\ \tilde{r}\to\infty$, (S.4)
where $\theta$ is the (macroscopic) contact angle at the wall. The Bond number $\mathrm{Bo}$ measures gravity vs. surface tension, and is related to the capillary length $\ell_{c}=\sqrt{\gamma/(\rho g)}$ via
$\mathrm{Bo}\;=\;\left(\frac{R}{\ell_{c}}\right)^{2}.$ (S.5)
[…]
$\tilde{H}^{\prime\prime}+\frac{1}{\tilde{r}}\,\tilde{H}^{\prime}-\mathrm{Bo}\,\tilde{H}=0,$ (S.6)
whose decaying solution is
$\tilde{H}(\tilde{r})\;\propto\;K_{0}\!\big{(}\lambda\,\tilde{r}\big{)},\qquad\lambda\equiv\sqrt{\mathrm{Bo}}\;=\;\frac{R}{\ell_{c}}.$ (S.7)
Asymptotically, $K_{0}(\lambda\tilde{r})\sim\sqrt{\pi/(2\lambda\tilde{r})}\,e^{-\lambda\tilde{r}}$, i.e. an exponential decay with a $1/\sqrt{\tilde{r}}$ prefactor. Thus, when lengths are scaled by $R$, the inverse decay length measured from the meniscus tail is precisely $\lambda=\sqrt{\mathrm{Bo}}$.
To impose the wall boundary condition $\tilde{H}^{\prime}(1)=-\cot\theta$ while retaining the linear Bessel form in the tail, we use
$\tilde{H}(\tilde{r})\;=\;C\;+\;\frac{\cot\theta}{\lambda\,K_{1}(\lambda)}\,K_{0}\!\big{(}\lambda\,\tilde{r}\big{)}$, (S.8)
where $C$ is a baseline offset that absorbs small imaging biases in the far field. Using $dK_{0}(x)/dx=-K_{1}(x)$, (S.8) yields
$\tilde{H}^{\prime}(1)=\frac{\cot\theta}{\lambda K_{1}(\lambda)}\bigl{(}-\lambda K_{1}(\lambda)\bigr{)}=-\cot\theta,$ (S.9)
as required. The fit therefore adjusts only $\lambda>0$ and $C$; the amplitude is fixed by the angle constraint. As soon as $\lambda$ is known, the fitted Bond number follows as
[…]
Laplace pressure from the fitted shape
The fitted profile can be used to compute the Laplace pressure jump $\Delta P$ along the interface. Two complementary approaches are used.
[…]
$\Delta\tilde{P}_{\text{lin}}(\tilde{r})\;\equiv\;\kappa_{\text{lin}}(\tilde{r})\;\approx\;-\,\mathrm{Bo}_{\text{fit}}\,\bigl{(}\tilde{H}(\tilde{r})-C\bigr{)}.$ (S.11)
The dimensional pressure follows from $\Delta P_{\text{lin}}=(\gamma/R)\,\Delta\tilde{P}_{\text{lin}}$.
Exact pressure (nonlinear curvature). Using the analytic derivatives of (S.8),
$\tilde{H}^{\prime}(\tilde{r})\;=\;-\,\frac{\cot\theta}{K_{1}(\lambda)}\,K_{1}\!\bigl{(}\lambda\tilde{r}\bigr{)},\qquad\tilde{H}^{\prime\prime}(\tilde{r})\;=\;\frac{\cot\theta\,\lambda}{2K_{1}(\lambda)}\Bigl{[}\,K_{0}\!\bigl{(}\lambda\tilde{r}\bigr{)}+K_{2}\!\bigl{(}\lambda\tilde{r}\bigr{)}\Bigr{]}$, (S.12)
[…]